GlueX TOF Meeting, August 16, 2016
GlueX Time-of-Flight Meeting
Tuesday, August 16, 2016
11:00 am EDT
JLab: CEBAF Center, Room F326/327
- Review of minutes from the previous meeting
- TOF Efficiency run10439 and run11508
- Run Plan
- Calibration Updates
- NIM paper
- The BlueJeans Meeting number is 350 531 998.
- Enter BlueJeans Meeting
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .
- FSU: Sasha Ostrovidov
- JLab: Fernando Barbosa, Brad Cannon, Paul Eugenio, Mark Ito (chair), Simon Taylor, Beni Zihlmann
There is a recording of this meeting on the BlueJeans site.
Review of minutes from the previous meeting
We looked over the minutes from August 9.
- Although it is expected that the timing is worse for cosmics runs than for beam data, we do not have a detailed understanding of the mechanism.
- Sasha reported a problem with the cosmic trigger: the energy threshold is expressed in 16 bits on the sum of all TOF counters and a single channel can cause the sum to exceed it. Beni will talk to Sasha about raising the threshold.
Beni showed scatter plots of efficiency for two runs: 10439 and 11508.
The plots show in the first column the distribution of track intersection points for tracks which match the TOF, for the vertical and horizontal paddles. The four plots in the second and third columns show, clockwise from top left, the charged track match efficiency for vertical paddles in 2-D bins of intersection position, the same for horizontal paddles, the efficiency for TOF points, and the efficiency in 2-D bins of charged track momentum versus polar angle.
Beni also showed efficiency as a function of run number for five paddle numbers both for horizontal and vertical plane.
Paddles 8 and 15 do not show much of an effect, but paddles 20 and 21 show unstable efficiencies near the end of the run. These are the ones closest to the beam. Paddle 5, although far from the beam, shows instability. This is not understood at present. The inefficiencies may be a combination of reduced gain at high rates and simply two-pulse resolution probability being reduced at high rate.
TDC Integral Non-Linearity Correction Calibration
Fernando reported that Ed Jastrzembski and Bryan Moffit are working on a scheme to obtain a non-linear correction for the CAEN TDCs. The TDCs are shipped from the factory with the correction pre-loaded, but that correction is based on a 40 MHz clock driving the boards; we run with 41.667 MHz. Turns out that the correction depends on the clock rate. The correction is different for each chip, but each chip services several channels.
The exact scheme for obtaining and downloading new calibration tables have not been worked out yet. The data will be obtained in situ in the Hall. We will need to schedule time for doing this before the run.
Paul will be turning to the NIM paper as one of his high priority tasks in the coming weeks, as well as looking into fixtures for gluing new tubes on the counters if that becomes necessary.